OpenStudy (anonymous):

is an=4n+16 equal to an=4+(n-1)16

OpenStudy (anonymous):

other options are an=16+(n-1)(4) an=16+4n an=20+(n)(4) an=20+(n-1)(4)

OpenStudy (precal):